# Neural reproducing kernel Banach spaces and representer theorems for deep networks

### v1

@@ 3. Neural RKBS
Next, let $\sigma:\mathbb{R}\to\mathbb{R}$ be a Lipschitz activation function such that $\sigma(0)=0$. This implies that
$\sigma(x)\in\ell^{2}(\mathbb{N})\qquad\text{for all }x\in\ell^{2}(\mathbb{N})\ ,$ (4)
@@
$\lim_{n\to\infty}\sigma(x_{n})=0\qquad\text{for all }x\in\ell^{2}(\mathbb{N})\ .$ (5)
Assuming $\sigma(0)=0$ is not restrictive. In fact, if $\sigma(0)\neq 0$, the choice of the activation function $\sigma^{\prime}=\sigma-\sigma(0)$ simply implies a scaling of the offsets. Then, we set
$\rho_{0}(x,n)=\begin{cases}1&n=0\\ x_{n}&n=1,\ldots,d\end{cases}\qquad x\in\mathbb{R}^{d}$
$\rho_{\ell}(x,n)=\begin{cases}1&n=0\\ \sigma(x_{n-1})&n\geq 1\end{cases}\quad x\in\ell^{2}(\mathbb{N})$ $\qquad\ell=1,\dots,L\ .$
We call the resulting (deep) integral RKBS $\mathcal{H}^{\text{deep}}$ a neural RKBS of depth $L$ and, if $L>1$, a deep neural RKBS.
Remark 3.1.
If $\sigma$ is Lipschitz with Lipschitz constant $C_{\sigma}$, then, for all $\ell=1,\dots,L$, $x,x^{\prime}\in\ell^{2}(\mathbb{N})$ and $n\in\mathbb{N}$,
$|\rho_{\ell}(x,n)-\rho_{\ell}(x^{\prime},n)|\leq|\sigma(x_{n-1})-\sigma(x^{\prime}_{n-1})|\leq C_{\sigma}|x_{n-1}-x^{\prime}_{n-1}|.$
General form of neural RKBS functions
An element $f^{\text{deep}}$ of the neural RKBS $\mathcal{H}^{\text{deep}}$ of depth $L$ is a composition of $L+1$ integral functions $f_{0}$,…,$f_{L}$,
@@
where each $f_{\ell}$ is defined via a measure $\mu_{\ell}$ by $f_{\ell}=f_{\mu_{\ell}}$. From layer 00 to layer $1$, we have
@@
$\quad\mu_{0}=\sum_{m=0}^{d}w^{(1)}_{m}\delta_{m}$
for a $d+1$-family of vectors $w^{(1)}_{0},\ldots w^{(1)}_{d}\in\ell^{2}(\mathbb{N})$. Let $b^{(1)}=w^{(1)}_{0}\in\ell^{2}(\mathbb{N})$ and define the bounded operator
$W^{(1)}\in B(\mathbb{R}^{d},\ell^{2}(\mathbb{N}))\qquad W^{(1)}x=\sum_{m=1}^{d}w^{(1)}_{m}x_{m}\ ,$
@@
$f_{0}(x)_{n}=\langle x,w^{(1)}_{n\cdot}\rangle_{\mathbb{R}^{d}}+b^{(1)}_{n}\qquad n\in\mathbb{N}\ ,$
@@
For $\ell=1,\dots,L-1$, from layer $\ell$ to layer $\ell+1$ we have
@@
$\|\mu_{\ell}\|_{\text{TV}}=\sum_{m=0}^{\infty}\|w^{(\ell+1)}_{m}\|_{\ell^{2}(\mathbb{N})}<+\infty\ .$ (6)
@@
$W^{(\ell+1)}\in B(\ell^{2}(\mathbb{N}),\ell^{2}(\mathbb{N}))\qquad W^{(\ell+1)}x=\sum_{m=1}^{+\infty}w^{(\ell+1)}_{m}x_{m-1}\ ,$
@@
$f_{\ell}(x)_{n}=\langle\sigma(x),w^{(\ell+1)}_{n\cdot}\rangle_{\ell^{2}}+b^{(\ell+1)}_{n}\qquad n\in\mathbb{N}\ ,$
@@
Finally, from layer $L$ to layer $L+1$, we have
@@
$W^{(L+1)}\in B(\ell^{2}(\mathbb{N}),\mathbb{R}^{p})\qquad W^{(L+1)}x=\sum_{m=1}^{+\infty}w^{(L+1)}_{m}x_{m-1}\ ,$
@@
$f_{L}(x)_{n}=\langle\sigma(x),w^{(L+1)}_{n\cdot}\rangle_{\ell^{2}}+b^{(L+1)}_{n}\qquad n=1,\ldots,p\ ,$ (8)
@@
We call the neural RKBS function $f$ an infinite-width neural network. Note that $\mathbb{N}$ parametrizes the width of the $L$ hidden layers $\ell=1,\dots,L$, while input and output layers $\ell=0$ and $\ell=L+1$ have fixed finite widths $d$ and $p$, respectively. Also note that infinite-width neural networks generalize finite-width neural networks, where the inner layers are generated by infinite-rank operators.
@@
Comparing our construction to the one proposed in [18], we remark that our networks are completely nonparametric, in the sense that every hidden layer has infinte width. In contrast, the networks in [18] are compositions of vector-valued shallow networks hence they have intermediate linear layers of finite width, corresponding to the finite output dimension of each shallow network. Such an architecture can be reinterpreted as an infinite-width network with finite-rank constraints. While imposing such constraints may enforce in practice some sort of beneficial stability, how to set or bound the ranks is an arbitrary choice.
Finite form of neural RKBS functions
We now show that the neural functions defined in 2.1 are correspond to measures $\mu_{1},\ldots,\mu_{L}$ having finite support. Indeed, under this assumption, for each $\ell=1,\ldots,L$,
$\mu_{\ell}=b^{(\ell+1)}\delta_{0}+\sum_{k=1}^{d_{\ell}}w^{(\ell+1)}_{k}\delta_{m^{(\ell)}_{k}}$
for some family of location $m^{(\ell)}_{1},\ldots,m^{(\ell)}_{d_{\ell}}\in\mathbb{N}\setminus\{0\}$ and some family of vectors $w^{(\ell+1)}_{1},\ldots,w^{(\ell+1)}_{k}\in\ell_{2}(\mathbb{N})$ if $\ell<L$ ($w^{(L+1)}_{1},\ldots,w^{(L+1)}_{k}\in\mathbb{R}^{p}$ if $\ell=L$). We define $f_{\ell}$ starting from the last layer. Since the support of $\mu_{L}$ is $\{0,m^{{L}}_{1},\ldots,m^{(L)}_{d_{L}}\}$, by (8) $f_{L}$ depends only on the variables $m^{(L)}_{1},\ldots,m^{(L)}_{d_{L}}$, so that we can regard $f_{L}$ as a function from $\mathbb{R}^{d_{L}}$ to $\mathbb{R}^{p}$ given by
@@
$W^{(L+1)}_{nk}=(w^{(L+1)}_{m_{k}^{(L)}})_{n-1}\qquad n=1,\ldots,p,\ k=1,\ldots,d_{L}\ .$
Since $f_{L}$ is defined on $\mathbb{R}^{d_{L}}$, regarded as finite dimensional subspace of $\ell_{2}(\mathbb{N})$, and denoted by $P:\ell_{2}(\mathbb{N})\to\mathbb{R}^{d_{L}}$ the corresponding projection
@@
$\mu_{\ell}\in\mathcal{M}(\{0,\ldots,d_{\ell}\},\mathbb{R}^{d_{\ell+1}})$
for some family $d_{0},d_{1},\dots,d_{L},d_{L+1}\in\mathbb{N}$ with $d_{0}=d$ and $d_{L+1}=p$. This means that
@@
Hence, for all $\ell=1,...,L+1$, let $W^{(\ell)}$ be the $n_{d_{\ell-1}}\times n_{d_{\ell}}$ matrix
@@
$f_{\ell}(x)=\begin{cases}W^{(1)}x+b^{(1)}&\ell=0\\ W^{(\ell+1)}\sigma(x)+b^{(\ell+1)}&\ell>1\ ,\end{cases}$
$f^{\text{deep}}=f_{L}\circ\ldots\circ f_{0}$
is a neural deep function according to 2.1, and we can rewrite $f^{\text{deep}}$ as
@@
that is, $f^{\text{deep}}$ is a neural network of depth $L$ and (finite) widths $d_{1},\dots,d_{L}$.
4. Representer theorems
In this section, we state and prove representer theorems for deep and neural RKBS, showing that neural networks can be seen as parametric solutions of a nonparametric model. We start by briefly recalling a basic supervised learning setting.
Let $\mathcal{X}$ be a set and $\mathcal{Y}$ a Banach space, called input and output space, respectively. Consider a Banach space $\mathcal{H}$ of functions $f:\mathcal{X}\to\mathcal{Y}$, called hypothesis space, and a loss function $\mathcal{L}:\mathcal{Y}\times\mathcal{Y}\to[0,\infty)$. Given $N$ samples
$(x_{i},y_{i})\in\mathcal{X}\times\mathcal{Y}\qquad i=1,\dots,N\ ,$
we want to solve the regularized empirical risk minimization problem
$\inf_{f\in\mathcal{H}}\mathcal{R}(f)+\|f\|_{\mathcal{H}}\ ,$ (9)
where
$\mathcal{R}(f)=\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}(f(x_{i}),y_{i})\ ,$
is the empirical error associated to the loss function $\mathcal{L}$ and the training points $(x_{i},y_{i})\in\mathcal{X}\times\mathcal{Y},i=1,\dots,N$.
To prove a representer theorem when $\mathcal{H}$ is a deep integral RKBS, we will need the following two fundamental assumptions on the basis functions $\rho_{\ell}$. We will see that both assumptions are easily satisfied by neural RKBS.
Assumption 1 (Pointwise form).
There is $\widetilde{\rho}_{\ell}:\mathbb{R}\times\Theta_{\ell}\to\mathbb{R}$ such that, for all $x\in\mathcal{X}_{\ell}$ and $\theta\in\Theta_{\ell}$,
$\rho_{\ell}(x,\theta)=\widetilde{\rho}_{\ell}(x(\theta),\theta)\ .$
Assumption 2 (Lipschitz condition).
There are $C_{\ell}>0$, $g_{\ell}\in\mathcal{C}_{b}(\Theta_{\ell},\mathcal{X}_{\ell})$ and $\beta_{\ell}\in\mathcal{C}_{0}(\Theta_{\ell})$ such that, for all $x,x^{\prime}\in\mathcal{X}_{\ell}$ and $\theta\in\Theta_{\ell}$,
$|\rho_{\ell}(x,\theta)-\rho_{\ell}(x^{\prime},\theta)|\leq C_{\ell}|\langle x-x^{\prime},g_{\ell}(\theta)\rangle_{\mathcal{X}_{\ell}}||\beta_{\ell}(\theta)|\ .$
For neural RKBS with activation function $\sigma$, 1 is satisfied with $\widetilde{\rho}_{0}(t,n)=t$ and $\widetilde{\rho}_{\ell}(t,n)=\sigma(t)$ for all $\ell=1,\dots,L$. Furthermore, by Remark 3.1, 2 is satisfied with $C_{\ell}=C_{\sigma}$ and $g_{\ell}(n)=\delta_{n-1}$ for all $\ell=1,\dots,L$.
Theorem 4.1 (Representer theorem for deep integral RKBS).
Let $\mathcal{H}$ be an integral RKBS of depth $L$ on $\mathcal{X}$ with values in $\mathcal{Y}$. Assume that $\mathcal{X}_{\ell+1}$, $\ell=0,\dots,L$, is Hilbert, $\mathcal{Y}=\mathcal{X}_{L+1}$ has finite dimension $d_{L+1}$, and $\mathcal{L}$ is continuous in the first entry. Moreover, assume that $\rho_{\ell}$ satisfies 1 for all $\ell=0,\dots,L$ and 2 for all $\ell=1,\dots,L$. Then, there are $d_{1},\dots,d_{L}<\infty$, $\theta^{(\ell)}\in\Theta_{\ell}^{d_{\ell}}$ and $w^{(\ell)}\in\mathcal{X}_{\ell}^{d_{\ell-1}}$ such that $f(x)=x^{(L+1)}$ defined by
$\begin{cases}x^{(0)}=x\\[5.0pt] x^{(\ell+1)}=\sum_{k=1}^{d_{\ell}}w^{(\ell+1)}_{k}\rho_{\ell}(x^{(\ell)},\theta^{(\ell)}_{k})\qquad\ell=0,\dots,L\end{cases}$
is a solution to problem (9). Moreover, for every $\ell=1,\dots,L$, $d_{\ell}\leq Nd_{\ell+1}$, and
$\|f\|_{\mathcal{H}}\leq\sum_{\ell=0}^{L}\sum_{k=1}^{d_{\ell}}\|w^{(\ell+1)}_{k}\|_{\mathcal{X}_{\ell+1}}\ .$
@@ 4. Representer theorems
Recall that in (9) the evaluation of an element $f=f_{0}\otimes f_{L}\in\mathcal{H}$ at point $x\in\mathcal{X}$ is given by $f(x_{i})=f^{\text{deep}}(x)$ where $f^{\text{deep}}$ is the composition of $f_{0},\ldots,f_{L}$. Furthermore, by construction each $f_{\ell}\in\mathcal{H}_{\ell}$ is parameterised by some measure $\mu_{\ell}\in\mathcal{M}_{\ell}=\mathcal{M}(\Theta_{\ell},\mathcal{X}_{\ell+1})$ according to (2). Taking into account account (3), the minimization problem (9) is equivalent to
$\inf_{\mu\in\mathcal{M}}\mathcal{R}(f_{\mu}^{\text{deep}})+\sum_{\ell=0}^{L}\|\mu_{\ell}\|_{\operatorname{TV}}=:\inf_{\mu\in\mathcal{M}}\mathcal{S}(\mu),$ (10)
where if $\mu=\mu_{0}\otimes\mu_{L}$, $f_{\mu}^{\text{deep}}$ is the composition of $f_{\mu_{0}},\ldots,f_{\mu_{L}}$.
By a classical argument, we first show that Problem (10) admits a solution. Pick a fixed $\nu\in\mathcal{M}$, and let $R=\mathcal{S}(\nu)$. Then Problem (10) is equivalent to
Indeed, take $\mu$ outside $\prod_{\ell=0}^{L}B_{\mathcal{M}_{\ell}}(R)$, then for some $\ell=0,\ldots,L$, $\|\mu_{\ell}\|_{\text{TV}}>R$, so that
which proves the equivalence. We now prove the existence of a minimizer of (11). By A.7, for all $x\in\mathcal{X}$ the map
$(\mu_{0},\dots,\mu_{L})\mapsto f_{\mu_{L}}\circ\cdots\circ f_{\mu_{0}}(x)$
is jointly continuous from $\prod_{\ell=0}^{L}B_{\mathcal{M}_{\ell}}(R)$, endowed with product topology induced by the weak-* topology of each $B_{\mathcal{M}_{\ell}}(R)$ to $\mathcal{Y}$. Since the $\mu_{\ell}\to\|\mu\|_{\operatorname{TV}}$ is weak-* continuous, the map $\mu\mapsto\mathcal{S}(\mu)$ is also continuous. Moreover, thanks to the Banach–Alaoglu theorem, the product $\prod_{\ell=0}^{L}B_{\mathcal{M}_{\ell}}(R)$ is $*{}^{*}$ compact. Hence, by the extreme value theorem, the problem (11) has at least a minimizer.
Let $\mu^{*}$ be any such solution. Denote $x_{i}^{(0)}=x_{i}$ and $x_{i}^{(\ell+1)}=f_{\mu^{*}_{\ell}}(x_{i}^{(\ell)})$ for $\ell=0,\dots,L$. Then, in view of A.4, a solution to (11) can be found by solving the following interpolation problems for all $\ell=0,\dots,L$:
Let us start from $\ell=L$. We want to apply A.2. To this end, let $\mathcal{U}=\mathcal{M}_{L}$ endowed with the $*{}^{*}$ topology. We define $\mathcal{A}:\mathcal{U}\to\mathbb{R}^{N\times d_{L+1}}$ by
$\mathcal{A}\mu=[f_{\mu_{L}}(x_{i})]_{i=1,\dots,N}\ .$
Then $\mathcal{A}$ is a surjective continuous linear operator from $\mathcal{U}$ onto $H=\operatorname{Ran}\mathcal{A}$, with $\dim(H)\leq Nd_{L+1}$. Moreover, the norm $G=\|\cdot\|_{\operatorname{TV}}$ is coercive on $\mathcal{U}$. Indeed, by the Banach–Alaoglu theorem, the balls $B_{\mathcal{M}_{L}}(r)$ are $*{}^{*}$ compact for every $r>0$. We define $F:H\to[0,\infty]$ by
$F(h)=\begin{cases}0&h_{i}=y_{i}\text{ for all $i=1,\dots,N$}\\ \infty&h_{i}\neq y_{i}\text{ for some $i=1,\dots,N$}\ .\end{cases}$
The function $F$ is the indicator function associeated to the singleton $\{(y_{0},\ldots,y_{L})\}$, so that it is convex, coercive and lower semi-continuous. Therefore, we can apply A.2, whence we obtain that (12) has a solution of the form
$\widetilde{\mu}_{L}=\sum_{k=1}^{d_{L}}c_{k}u_{k}$
for some $d_{L}\leq Nd_{L+1}$, $c_{k}>0$ and $u_{k}\in\operatorname{Ext}(B_{\mathcal{M}_{L}}(1))$. By A.3, for each $k$ there are $y_{k}\in\operatorname{Ext}(B_{\mathcal{Y}}(1))$ and $\theta^{(L)}_{k}\in\Theta_{L}$ such that
$u_{k}=y_{k}\cdot\delta_{\theta^{(L)}_{k}}\ .$
Thus, defining $w^{(L+1)}_{k}=c_{k}y_{k}$, we get
$\widetilde{\mu}_{L}=\sum_{k=1}^{d_{L}}w^{(L+1)}_{k}\delta_{\theta^{(L)}_{k}}\ .$
We now come to $\ell=L-1$. Thanks to 1, we can regard $f_{\widetilde{\mu}_{L}}$ as defined on
$\widetilde{\mathcal{X}}_{L}=\mathbb{R}^{\operatorname{{supp}}(\widetilde{\mu}_{L})}\ ,$
and restrict (12) to
$\inf_{\mu_{L-1}\in\mathcal{M}(\Theta_{L-1},\widetilde{\mathcal{X}}_{L})}\|\mu_{L-1}\|_{\operatorname{TV}}\quad\text{subject to}\quad f_{\mu_{L-1}}(x_{i}^{(L-1)})={x_{i}^{(L)}}_{|\operatorname{{supp}}(\widetilde{\mu}_{L})}\quad i=1,\dots,N\ .$
We use again A.2, together with A.3, as in the previous step ($\ell=L$). In particular, this time we have $\dim(H)\leq Nd_{L}$. Thus, we find a solution of the form
for some $d_{L-1}\leq Nd_{L}$.
Iterating the argument for $\ell=L-2,\dots,0$, the main claim follows. The bound on $d_{\ell}$ is also clear by iteration. For the bound on $\|f\|_{\mathcal{H}}$, by definition we have
$\|f\|_{\mathcal{H}}=\sum_{\ell=0}^{L}\|f_{\ell}\|_{\mathcal{H}_{\ell}}=\sum_{\ell=0}^{L}\|f_{\widetilde{\mu}_{\ell}}\|_{\mathcal{H}_{\ell}}\leq\sum_{\ell=0}^{L}\|\widetilde{\mu}_{\ell}\|_{\operatorname{TV}}\ ,$
where, once again by A.2,
$\|\widetilde{\mu}_{\ell}\|_{\operatorname{TV}}=\sum_{k=1}^{d_{\ell}}c_{k}\ .$
But since $y_{k}\in\operatorname{Ext}(B_{\mathcal{X}_{\ell+1}}(1))$, we have $\|y_{k}\|_{\mathcal{X}_{\ell+1}}=1$, whence
which concludes the proof. ∎
In the special case of a neural RKBS, 4.1 implies the following result.
Corollary 4.2 (Representer theorem for neural RKBS).
Let $\mathcal{H}$ be a neural RKBS of depth $L$ on $\mathbb{R}^{d}$ with valued in $\mathbb{R}^{d_{L+1}}$. Assume that $\mathcal{L}$ is continuous in the first entry. Then, there are $d_{1},\dots,d_{L}<\infty$, $W^{(\ell)}\in\mathbb{R}^{d_{\ell}\times d_{\ell-1}}$ and $b^{(\ell)}\in\mathbb{R}^{d_{\ell}}$ such that $f(x)=x^{(L+1)}$ defined by
$\begin{cases}x^{(1)}=W^{(1)}x\\[5.0pt] x^{(\ell+1)}=W^{(\ell+1)}\sigma(x^{(\ell)}-b^{(\ell)})\qquad\ell=1,\dots,L\end{cases}$
is a solution to problem (9). Moreover, for every $\ell=1,\dots,L$, $d_{\ell}\leq Nd_{\ell+1}$, and
$\|f\|_{\mathcal{H}}\leq\sum_{\ell=0}^{L}\sum_{k=1}^{d_{\ell}}\biggl{(}\sum_{j=1}^{d_{\ell+1}}|W^{(\ell+1)}_{jk}|^{2}\biggr{)}^{1/2}\ .$
@@
Our result shows that deep neural networks are solutions to regularized empirical risk minimization over neural RKBS, thus justifying neural RKBS as hypothesis spaces for deep learning. Moreover, it provides an upper bound on the network width depending on sample size and input/output dimensions. Finally, it shows that the regularization norm is controlled by the $\ell^{1}$ norm of the $\ell^{2}$ norms of the weights of the network. Finer characterizations of the Banach structure can be obtained using the specific form of the activation function and the functional properties that this induces (see [18] for the ReLU).
5. Conclusion

### v2

@@ 6. Discrete Neural RKBS
and let $\sigma:\mathbb{R}\to\mathbb{R}$ be a Lipschitz activation function such that $\sigma(0)=0$. Then, set
$\rho_{0}(x,n)=\begin{cases}1&\hskip 29.0ptn=0\\ x_{n}&\hskip 29.0ptn=1,\ldots,d\end{cases}\quad,\qquad$ $x\in\mathbb{R}^{d}$
$\rho_{\ell}(x,n)=\begin{cases}1&n=0\\ \sigma(x_{n-1})&n\geq 1\end{cases}\quad,$ $x\in\ell^{2}(\mathbb{N}),\qquad\ell=1,\dots,L\ .$
We call the resulting space $\mathcal{H}^{\rm deep}$ a discrete neural RKBS.
Remark 6.2.
Some observations on Definition 6.1 are in order. First, note that $\mathcal{X}_{0}$ can be thought of as a function space on $\Theta_{0}$ by putting $x(0)=1$ for all $x\in\mathcal{X}_{0}$. Second, the Lipschitzianity of $\sigma$ implies that
$\sigma(x)\in\ell^{2}(\mathbb{N}),\qquad\text{for all }x\in\ell^{2}(\mathbb{N}),$ (22)
@@
$\lim_{n\to\infty}\sigma(x_{n})=0,\qquad\text{for all }x\in\ell^{2}(\mathbb{N})\ .$ (23)
Finally, assuming $\sigma(0)=0$ is not restrictive. In fact, if $\sigma(0)\neq 0$, the choice of the activation function $\sigma^{\prime}=\sigma-\sigma(0)$ simply implies a scaling of the offsets.
General form of discrete neural RKBS functions
An element $f^{\rm deep}$ of the neural RKBS $\mathcal{H}^{\rm deep}$ is a composition of $L+1$ integral functions $f_{0}$$f_{L}$, where each $f_{\ell}$ is defined via a measure $\mu_{\ell}$ by $f_{\ell}=f_{\mu_{\ell}}$. We now derive an explicit expression for elements in a discrete neural RKBS, showing that these functions correspond to infinite-width neural networks with a countably infinite number of neurons per hidden layer, see Figure 2.
@@
[…]
Figure 2. Architecture of discrete neural RKBS functions.
Case $\ell=0$. From layer 0 to layer $1$, we have
@@
$\quad\mu_{0}=\sum_{m=0}^{d}w^{(1)}_{m}\delta_{m},$
for a family of $d+1$ vectors $w^{(1)}_{0},\ldots w^{(1)}_{d}\in\ell^{2}(\mathbb{N})$. Let $b^{(1)}=w^{(1)}_{0}\in\ell^{2}(\mathbb{N})$ and define the bounded operator
$W^{(1)}\in B(\mathbb{R}^{d},\ell^{2}(\mathbb{N})),\qquad W^{(1)}x=\sum_{m=1}^{d}w^{(1)}_{m}x_{m}.$
@@
$f_{0}(x)_{n}=\langle x,w^{(1)}_{n\cdot}\rangle_{\mathbb{R}^{d}}+b^{(1)}_{n},\qquad n\in\mathbb{N},$
@@
Case $\ell=1,\dots,L-1$. From layer $\ell$ to layer $\ell+1$ we have
@@
$\|\mu_{\ell}\|_{\text{TV}}=\sum_{m=0}^{\infty}\|w^{(\ell+1)}_{m}\|_{\ell^{2}(\mathbb{N})}<\infty.$ (24)
@@
$W^{(\ell+1)}\in B(\ell^{2}(\mathbb{N}),\ell^{2}(\mathbb{N})),\qquad W^{(\ell+1)}x=\sum_{m=1}^{+\infty}w^{(\ell+1)}_{m}x_{m-1},$
@@
$f_{\ell}(x)_{n}=\langle\sigma(x),w^{(\ell+1)}_{n\cdot}\rangle_{\ell^{2}}+b^{(\ell+1)}_{n},\qquad n\in\mathbb{N},$
@@
Case $\ell=L$. Finally, from layer $L$ to layer $L+1$, we have
@@
$W^{(L+1)}\in B(\ell^{2}(\mathbb{N}),\mathbb{R}^{p}),\qquad W^{(L+1)}x=\sum_{m=1}^{+\infty}w^{(L+1)}_{m}x_{m-1},$
@@
$f_{L}(x)_{n}=\langle\sigma(x),w^{(L+1)}_{n\cdot}\rangle_{\ell^{2}}+b^{(L+1)}_{n},\qquad n=1,\ldots,p,$ (26)
@@
We stress that the width of the $L$ hidden layers $\ell=1,\dots,L$ is infinite and countable (the neurons are parameterized by $\mathbb{N}$), while input and output layers $\ell=0$ and $\ell=L+1$ have fixed finite widths $d$ and $p$, respectively. Also note that infinite-width neural networks generalize finite-width neural networks, where the inner layers are generated by infinite-rank operators.
@@
Comparing our construction to the one proposed in [22], we remark that our networks do not have any rank constraint, in the sense that every hidden layer has infinte width.
Finite form of discrete neural RKBS functions
We now show that the neural functions defined in 2.1 correspond to measures $\mu_{1},\ldots,\mu_{L}$ having finite support. Indeed, under this assumption, for each $\ell=1,\ldots,L$,
$\mu_{\ell}=b^{(\ell+1)}\delta_{0}+\sum_{k=1}^{d_{\ell}}w^{(\ell+1)}_{k}\delta_{m^{(\ell)}_{k}}$,
for some $m^{(\ell)}_{1},\ldots,m^{(\ell)}_{d_{\ell}}\in\mathbb{N}\setminus\{0\}$ and some $w^{(\ell+1)}_{1},\ldots,w^{(\ell+1)}_{k}$ that are in $\ell_{2}(\mathbb{N})$ if $\ell<L$, and in $\mathbb{R}^{p}$ if $\ell=L$.
We define $f_{\ell}$ starting from the last layer. Since the support of $\mu_{L}$ is $\{0,m^{(L)}_{1},\ldots,m^{(L)}_{d_{L}}\}$, by (26) $f_{L}$ depends only on the variables $m^{(L)}_{1},\ldots,m^{(L)}_{d_{L}}$, so that we can regard $f_{L}$ as a function from $\mathbb{R}^{d_{L}}$ to $\mathbb{R}^{p}$ given by
@@
$W^{(L+1)}_{nk}=(w^{(L+1)}_{m_{k}^{(L)}})_{n-1},\qquad n=1,\ldots,p,\qquad k=1,\ldots,d_{L}.$
Since $f_{L}$ is defined on $\mathbb{R}^{d_{L}}$, regarded as a finite-dimensional subspace of $\ell_{2}(\mathbb{N})$, denoting by $P:\ell_{2}(\mathbb{N})\to\mathbb{R}^{d_{L}}$ the corresponding projection
@@
$\mu_{\ell}\in\mathcal{M}(\{0,\ldots,d_{\ell}\},\mathbb{R}^{d_{\ell+1}}),$
for some $d_{0},d_{1},\dots,d_{L},d_{L+1}\in\mathbb{N}$ (with $d_{0}=d$ and $d_{L+1}=p$). This means that
@@
For all $\ell=1,\dots,L+1$, let $W^{(\ell)}$ be the $n_{d_{\ell-1}}\times n_{d_{\ell}}$ matrix
@@
$f_{\ell}(x)=\begin{cases}W^{(1)}x+b^{(1)}&\ell=0\\ W^{(\ell+1)}\sigma(x)+b^{(\ell+1)}&\ell>1\end{cases},$
$f^{\rm deep}=f_{L}\circ\ldots\circ f_{0},$
is a neural deep function according to Definition˜2.1, and we can rewrite $f^{\rm deep}$ as
@@
that is, $f^{\rm deep}$ is a neural network of depth $L$ and (finite) widths $d_{1},\dots,d_{L}$.
Remark 6.3.
For discrete neural RKBS, Equation˜10 is satisfied replacing the basis functions in Definition 6.1 by $\rho_{\ell}(x,n)=\sigma(x_{n-1})\beta_{n-1}$ for all $\ell=1,\dots,L$, with $\sigma\colon\mathbb{R}\to\mathbb{R}$ Lipschitz and $\beta:\mathbb{N}\to\mathbb{R}$ a positive sequence converging to zero. Furthermore, Equation˜11 is satisfied with $C_{\ell}=C_{\sigma}$ and $g_{\ell}(n)=\delta_{n-1}$ for all $\ell=1,\dots,L$. Indeed, for all $x,x^{\prime}\in\ell^{2}(\mathbb{N})$ and $n\in\mathbb{N}$,
$|\rho_{\ell}(x,n)-\rho_{\ell}(x^{\prime},n)|\leq|\sigma(x_{n-1})-\sigma(x^{\prime}_{n-1})||\beta_{n-1}|\leq C_{\sigma}|x_{n-1}-x^{\prime}_{n-1}||\beta_{n-1}|.$
Corollary 6.4 (Representer theorem for discrete neural RKBS).
Under the assumptions of ˜5.3, if $\mathcal{H}^{\rm deep}$ is a discrete neural RKBS, then the claim of ˜5.5 holds true. Moreover,
$\Phi(f^{\rm deep})\leq\sum_{\ell=0}^{L}\sum_{k=1}^{d_{\ell}}\biggl{(}\sum_{j=1}^{d_{\ell+1}}|W^{(\ell+1)}_{jk}\beta_{k}^{-1}|^{2}\biggr{)}^{1/2}.$
@@
Since $\Theta_{0}=\{0,\ldots,d\}$, without loss of generality we can assume $d_{0}=d+1$ and $\theta_{n}^{(0)}=n$ for all $k=0,\ldots,d_{0}$. Taking into account that $\rho_{0}(x,0)=1$ and $\rho_{0}(x,k)=x_{k}$ if $1\leq k\leq d$, we have
$x^{(0)}=(1,x)\in\mathbb{R}\times\mathbb{R}^{d}.$
Since, for $\ell=1,\ldots,L$, $\Theta_{\ell}=\mathbb{N}$ and $K_{\ell}(\cdot,n)=e_{n-1}$ where $\{e_{n}\}_{n\in\mathbb{N}}$ is the canonical base of $\mathcal{X}_{\ell}=\ell_{2}(\mathbb{N})$, up to a permutation, $V^{(\ell)}$ is the identity and the claim follows. ∎
@@
The above result shows that deep neural networks with finite width at each hidden layer are optimal, in the sense that they are solutions of empirical risk minimization over neural RKBS. Moreover, it provides an upper bound on the network width depending on sample size and input/output dimensions. Finally, it shows that the regularization norm is controlled by the $\ell^{1}$ norm of the $\ell^{2}$ norms of the weights of the network.
7. Conclusions and future work